OpenStudy (anonymous):

[4.02]What is the y−coordinate of the y−intercept of the line that passes through the points (−4, −4) and (4, 8)?

OpenStudy (aripotta):

gotta find the slope first

OpenStudy (anonymous):

which iss...

OpenStudy (anonymous):

that...

OpenStudy (aripotta):

yea, that lol

OpenStudy (anonymous):

8-(-4)

OpenStudy (anonymous):

4-(-4)

OpenStudy (anonymous):

which turns to..

OpenStudy (anonymous):

2/3 right?

OpenStudy (anonymous):

12/8

OpenStudy (anonymous):

and yeah.

OpenStudy (anonymous):

actually....

OpenStudy (anonymous):

wait no 3/2

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

mhmm,

OpenStudy (anonymous):

and then y=mx+b

OpenStudy (anonymous):

so how do i find the y-intercept?

OpenStudy (anonymous):

is your formula.

OpenStudy (aripotta):

y = mb + b. use a coordinate and plug in its x and y. then put in the slope and solve for b

OpenStudy (anonymous):

so... it should be

OpenStudy (anonymous):

remember m is your slope X3

OpenStudy (anonymous):

Just a quick reminder.

OpenStudy (aripotta):

mx + b*

OpenStudy (anonymous):

so y = 3/2 + what?

OpenStudy (anonymous):

dont i have to put it in point-slope form?

OpenStudy (anonymous):

No, you don't....

OpenStudy (anonymous):

Point slope is different.

OpenStudy (aripotta):